# Strong Coupling Constant Determination from the new CTEQ-TEA Global QCD Analysis

### v1

@@ I Introduction
The strong coupling constant $\alpha_{s}$ is a fundamental parameter of the QCD Lagrangian that enters theoretical computations for all hadronic scattering processes. A precise and accurate determination of $\alpha_{s}(Q)$ and its dependence on the energy scale $Q$ is of vital importance for the current precision-frontier program at the Large Hadron Collider (LHC), where vast amounts of data are being gathered and push the boundary of precision measurements to the next level [1]. The upcoming high-luminosity phase of the LHC (HL-LHC) will further increase the volume of high-precision data, demanding a matching increase in both the detector capabilities and precision of theoretical calculations to properly validate relationships between fundamental parameters of the Standard Model (SM), constrain properties of the Higgs boson, and search for new physics interactions. For instance, the running of $\alpha_{s}$ at the highest achieved energies may reveal the presence of new colored particles and influence the scale at which the SM couplings may unify. Variations in the observed $\alpha_{s}(M_{Z})$ at the scales of order of the mass of $Z$ boson, $M_{Z}$, can affect projections for the range of energies at which the electroweak (EW) vacuum might shift from a metastable to a stable configuration.
@@
Hence, the QCD coupling and its estimated uncertainty play a pivotal dual role of characterizing the universal force of nature and at the same time serving as a precise theoretical input for many standard candle processes in ongoing experiments. In particular, top-quark pair and Higgs boson production cross sections are very sensitive to the assumed $\alpha_{s}(M_{Z})$: the associated uncertainty, entering at hadron colliders with the correlated PDF uncertainty, can potentially exceed those due to the scale dependence and experimental errors [2, 3].
@@
The Particle Data Group (PDG) currently reports the uncertainty on $\alpha_{s}(M_{Z})$ of about 0.8% based on an analysis of combined world experimental data, resulting in $\alpha_{s}(M_{Z})=0.118\pm 0.0009$  [4]. The corresponding relative uncertainty exceeds those achieved for the fundamental couplings of electroweak and gravitational interactions. The new twist is that the latest determinations of $\alpha_{s}$ from high-statistics experimental measurements and lattice QCD may individually quote a comparable, and soon a smaller, uncertainty than the PDG world-average uncertainty of $\pm 0.0009$. Along these lines, the ATLAS collaboration measured $\alpha_{s}(M_{Z})=0.1183\pm 0.0009$ [5] using data of $Z$ boson transverse-momentum distributions at $\sqrt{s}=8$ TeV. The CMS collaboration simultaneously extracted $\alpha_{s}(M_{Z})$ and PDFs [6] from a global QCD analysis at NNLO using double-differential cross section measurements of inclusive jet production at $\sqrt{s}=$ 2.76, 7, 8, and 13 TeV combined with deep inelastic data from HERA. They obtained $\alpha_{s}(M_{Z})=0.1176^{+0.0014}_{-0.0016}$. Other recent relevant $\alpha_{s}$ determinations from the CMS and ATLAS collaborations can be found in Refs. [7, 8, 5, 9]. Several determinations from deep inelastic scattering (DIS) at $ep$ collider HERA find a somewhat lower $\alpha_{s}$ value than at the LHC, such as $\alpha_{s}(M_{Z})=0.1142\pm 0.0028$ determined by H1 collaboration in a simultaneous fit of $\alpha_{s}$ and proton PDFs [10]. Independently, lattice QCD determines $\alpha_{s}$ from several types of nonperturbative computations. The Flavor Lattice Averaging Group (FLAG), systematizes and cross validates these determinations. It recently reports $\alpha_{s}(M_{Z})=0.1183\pm 0.0007$ [11].
@@
The studies of $\alpha_{s}$ thus enter a new phase when its accuracy can be potentially pushed to as low as 0.1% across diverse observations [1], while in the meantime novel challenges will need to be confronted both in the individual $\alpha_{s}$ measurements and their combination. It is expected that, at this precision level, the growing complexity of involved systematic factors raises the chance for misestimations. When the misestimations go unnoticed – an unwanted common occurence encountered in complex studies across many disciplines [12] – the uncertainty tends to be underestimated; some measurements of the parameter are not replicated by its independent determinations. When replicability degrades, a PDG or another combination may need to inflate the stated uncertainties of measurements in order to make them statistically compatible. As an illustration, the Committee on Data of the International Science Council inflates the errors of the highly precise measurements of the Newton’s gravitational constant $G$ by a factor of 3.9 in order to meaningfully combine them [13]. Within its small span, the combined uncertainty band on $G$ does not overlap significantly with those of several most precise measurements, meaning it does not replicate them. This indicates a general trend that the measurements of fundamental constants eventually experience the tradeoff between the precision and replicability, when raising the former depletes the confidence in the latter. How far can the precision measurements of $\alpha_{s}$ go while remaining replicable? What can be done to control the uncertainty estimates and postpone the onset of non-replicability?
@@
Our article will critically explore these questions on the case of the precision determination of the $\alpha_{s}(M_{Z})$ from the new CTEQ Tung Et Al. (CTEQ-TEA) global QCD analysis of nucleon scattering data, named “CT25” [14]. The primary goal of the global QCD analysis is to determine parametrizations of parton distribution functions (PDFs), as well as $\alpha_{s}(M_{Z})$ and other QCD parameters, from a simultaneous fit of a large, diverse collection of experimental measurements. The methodology of the global analysis offers the most complete look into the correlated constraints on $\alpha_{s}$, heavy-quark masses, and the PDFs arising from a variety of scattering experiments that simultaneously depend on them. This completeness gives the global analysis an edge over the standalone $\alpha_{s}$ determinations from the nucleon data, which typically neglect at least some of the correlation with the PDFs or other experiments. As such, and given the growing dominance of hadron scattering processes in experimental $\alpha_{s}$ determinations, the global analysis of PDFs is well posed for tests of uncertainty quantification and replicability in the $\alpha_{s}$ studies. Many combined analyses of the PDFs and $\alpha_{s}$ have been performed by our CTEQ-TEA group [15, 16] and other groups [17, 18, 19, 20, 21, 22] over the years. The strong coupling $\alpha_{s}(M_{Z})$ has recently been extracted from global QCD analyses, employing an approximate PDF evolution at the next-to-next-to-next-to-leading order (aN3LO) [23] in QCD by the MSHT group, and from an analysis using an aN3LO QCD evolution combined with the photon PDF at the next-to-leading order (NLO) in QED [24] by the NNPDF collaboration.
@@
The CT25 global analysis determines $\alpha_{s}$ at the next-to-next-to-leading order (NNLO) in QCD by employing an improved methodology as well as an extended global data set in which precision LHC Run-2 measurements with high integrated luminosity augment those included in the CT18 study [16]. The novel feature of our study is that we determine the central value and uncertainty of $\alpha_{s}(M_{Z})$ by several methods to obviate the dependence on the correlation with PDFs, selection of experiments, and systematic uncertainties in the experiments.
@@
The rest of this manuscript is organized as follows. In Sec. II, we briefly describe the methodologies and advancements in the CT25 global PDF analysis. We summarize the new data sets included in the CT25 baseline, definition of the figure-of-merit function $\chi^{2}$, and implementation of correlated systematic errors provided by the experiments. Section III.1 explores the impact of the choice of PDF parametrizations on the $\alpha_{s}$ determination. We also review how $\alpha_{s}(M_{Z})$ can be determined from the $\chi^{2}$ scan over $\alpha_{s}(M_{Z})$, and which criteria are applicable for determination of the $\alpha_{s}(M_{Z})$ and its error in this methodology.
@@
Sec. III.2 critically examines the impact of the implementation of experimental systematic uncertainties on the extracted value of $\alpha_{s}(M_{Z})$, and it discusses handling of some observed disagreements among data sets in the baseline. Section IV is dedicated to the exploration of uncertainty prescriptions (tolerances) according to the commonly used global and dynamical implementations. For the latter definition, we examine how the user-specified combination of published data sets into clusters modifies the quantitative error estimates. We point out that either the dynamical tolerance in the Hessian method or $K$-folding in the Monte-Carlo method may produce uncertainties that depend on user-chosen clustering of data sets. We introduce the concept of data clustering safety that must be addressed in such situations.
@@
Sec. V describes Bayesian statistical models utilized for the combination of the results for $\alpha_{s}(M_{Z})$ and its uncertainties, while Sec. VI obtains and compares the uncertainties using a variety of methods. This Section quotes our final $\alpha_{s}$ estimate. Sec. VII contains our conclusions.
New experiments in the CT25 global fit
@@
Inclusive jet production
553 ATLAS 8 TeV IncJet   [36] 171 $1.60^{+0.10}_{-0.06}$ ( $1.57^{+0.12}_{-0.07}$ )
554 ATLAS 13 TeV IncJet  [37] 177 $1.32^{+0.06}_{-0.05}$ ( $1.26^{+0.09}_{-0.03}$ )
555 CMS 13 TeV IncJet   [38] 78 $1.10^{+0.10}_{-0.05}$ ( $1.12^{+0.17}_{-0.13}$ )
@@
CT25prel, with augmented CT18A data set 4450 $1.20\pm 0.010$ $\approx 0.118$
@@
Table 1: The upper part of the table lists the newly added data sets in the CT25 analysis, contributed by measurements in lepton-pair, top-quark pair, and single-inclusive jet production from the LHC Runs 1 and 2. They contain $N_{\textrm{pt}}=769$ new data points in total. For these experiments, we report the $\chi^{2}/N_{\textrm{pt}}$ for the default CT25 and prior CT18A PDF ensembles, assuming $\alpha_{s}(M_{Z})=0.118$. Section II.2 provides the definition of $\chi^{2}$. The errors are asymmetric PDF uncertainties computed with a master formula in Ref. [39], and with the error PDFs defined at 68% CL according to the two-tier PDF tolerance employed in CT18. The lower part of the table shows the total numbers of points, $\chi^{2}/N_{\textrm{pt}}$, and approximate preferred $\alpha_{s}(M_{Z})$ in the default and preliminary CT25 fits described in the text, as well as in the CT18A fit.
II Relevant aspects of the CT25 global analysis
II.1 Selection of data sets
As many general-purpose CTEQ-TEA fits, the CT25 NNLO global analysis took place over several years and now culminates with the release of the official CT25 PDF ensembles that update the previous generation of the PDFs [16] consisting of the default CT18 NNLO error set and three alternative, yet equally acceptable PDF ensembles, CT18Z, A, and X. A significant part of this effort went into an exploration of a large number of new candidate measurements for the extension of the CT18 baseline data set, focusing primarily on high-precision production of lepton pairs, hadronic jets, and top quark-antiquark pairs in the LHC Runs 1 and 2. From many such examined data sets, we selected several precise and mutually consistent ones that we included into the final CT25 release. We documented these preliminary explorations in three publications [40, 41, 42].
These newly added LHC data sets in CT25 are listed in the upper part of Table 1 and reviewed in Appendix A. As its data baseline, CT25 started with the CT18A [16] global data set, which already included a combination of measurements in deep-inelastic scattering and production of lepton pairs, hadronic jets, and top-quark pairs at both fixed-target and collider energies. Notably, the CT18A baseline included a precision data set on $W$ and $Z$ production in the ATLAS experiment at 7 TeV [43], which offered the first evidence for the enhanced strangeness PDF at $x\approx 0.02$, in some disagreement with charged-current DIS experiments. The even more recent LHC measurements included in the CT25 fit are consistent with this strangeness enhancement.
The inclusion of the new LHC measurements, added on top of the CT18A baseline without replacing or discarding any of the previously fitted experiments, increased the number of data points, $N_{\textrm{pt}}$, by 769, the majority of which were from inclusive jet production and Drell-Yan triple-differential $Z$-boson production at ATLAS, cf. Table 1. This version of the CT25 fit, which we will call “CT25prel”, has the largest $N_{\textrm{pt}}=4450$ among the various CT25 candidate fits. This data set is not unique, however: other well-motivated data combinations are possible, as was also in the CT18 case. When all data sets are retained in the CT25 baseline, we observe a reasonable, although not ideal, agreement among the global data, with significant pulls on the PDFs from the newly-fitted data. The most precise new experiments in the CT25 fits tend to exhibit an increase in $\chi^{2}/N_{\textrm{pt}}$ in a pattern similar to that observed in the CT18 global analysis. The new Drell-Yan and $t\bar{t}$ production data consistently prefer a softer gluon at intermediate-to-large $x$ ($0.05\leq x\leq 0.7$), while jet production data pull the gluon in the opposite direction in the same kinematic range. Overall, the cumulative effects of these measurements, when they are combined, amounts to having a softer gluon at the intermediate-to-large $x$, as compared to CT18A.
To representatively investigate the range of PDF solutions in this new setting, we explored other descendant, also acceptable, branches of the CT25prel fit. The alternative choices made in these fits reflect the complexities present within the global fitting framework, many of which were highlighted in the preceding CT18 publication [16], as well as in a review [25]. [^1] In the context of CT18, we opted to publish four PDF ensembles to quantify some of these possibilities. In the end, we obtained the default CT25 fit, or just “CT25”, as follows. Compared to CT18A and CT25prel, we replaced the H1 measurements of charm and bottom semi-inclusive structure functions [44, 45] by their HERA combination [46], which moderately increased the corresponding $\chi^{2}$ without substantially modifying the PDFs. We added the E906/SeaQuest measurement of the $\sigma_{pd}/(2\sigma_{pp})$ ratio in fixed-target Drell-Yan pair production [47]. We removed four data sets (336 points) on inclusive charged-current DIS on an iron target from the CDHSW [48] and CCFR [49, 50] collaborations, as they are in some tension with each other and increase dependence of CT25 results on the nuclear modifications of proton PDFs. [^2] The default CT25 fit nevertheless retains CCFR and NuTeV data sets on dimuon production in CC DIS on the isoscalar target, as those provide unique constraints on the strangeness PDF at $x>0.05$. In both CT18 and CT25 fits, we apply a correction from [51] to approximate nuclear modifications in such experiments. The weakly sensitive determination of the longitudinal structure function $F_{L}(x_{B},Q^{2})$ by the H1 collaboration [52] was also omitted without a noticeable impact on the PDFs.
@@ II.1 Selection of data sets
Turning back to Table 1, the last column in its upper part reports the ratios the $\chi^{2}/N_{\rm pt}$ values for each listed experiment using the PDFs from the default CT25 NNLO and CT18A NNLO global fits (with the latter in parentheses). The errors express the corresponding asymmetric PDF uncertainties that are evaluated at the 68% credibility level (CL). [^3] The PDF fits base themselves on the Bayesian paradigm, and hence the interpretation of their errors on the PDFs and $\alpha_{s}$ as “credibility intervals” is appropriate. That said, these errors are also used to predict the frequency of future measurements using the determined parameters, in which case the users may view them as “confidence intervals”. Fitting the new experiments in the CT25 framework generally improves their $\chi^{2}$.
@@
The lower part of Table 1 compares the total numbers of points and $\chi^{2}/N_{\textrm{pt}}$ for the default CT25, CT25prel, and their CT18A predecessor. The last column of the lower part also shows the approximate preferred value of $\alpha_{s}(M_{Z})$ in these three fits, which is further adjusted by modifications discussed in the rest of the article.
@@
Upon the addition of the LHC data sets in CT25prel, the reference $\alpha_{s}(M_{Z})$ increased from a range 0.1164-0.1169 in the CT18 series to $\approx 0.118$. To a large extent, this change is due to the preference for a softer gluon at the intermediate-to-large $x$, which is compensated by some increase in $\alpha_{s}(M_{Z})$.
@@
In the default CT25 fit, the reference $\alpha_{s}(M_{Z})$ is further increased to $\approx 0.1185$ primarily because of the exclusion of the CDHSW inclusive charged-current DIS data set [48] on iron. The exclusion of the counterpart CCFR measurement had no appreciable effect. In the CT18 publication, we emphasized that the CDHSW data set is generally fitted well, yet its $Q$ dependence differs from that of an analogous CCFR observation [53], which in turn may reflect differences in calibration and resolution of the two detectors [54]. The Lagrange Multiplier scans in the context of CT18 (see its Fig. 21) showed that CDHSW prefers a far larger gluon $g(x,Q)$ at $x>0.1$ than the other experiments. This preference, in turn, suppresses the preferred $\alpha_{s}(M_{Z})$. For this reason, the two CDHSW data sets were included in CT18/CT18A/CT18X, but not in CT18Z. Either way, the impact of adding or removing the inclusive neutrino-iron DIS data sets is weak and comparable to other less known factors. The $\chi^{2}/N_{\textrm{pt}}$ is smaller when these CDHSW and CCFR experiments are included, so they do not manifestly disagree with the rest.
@@
Our general observation is then that the described revisions in the CT25 data set, as compared to the CT18A one, tend to elevate the preferred $\alpha_{s}(M_{Z})$ to above 0.118 for the default settings of the global fit. We should next emphasize that this trend may change considerably under reasonable modifications in the fit’s settings, reflecting in part inconsistencies (tensions) among the input data sets, and in part the modeling of experimental and theoretical systematic uncertainties.

### v2

@@ I Introduction
The strong coupling constant $\alpha_{s}$ is a fundamental parameter of the QCD Lagrangian that enters theoretical computations for all hadronic scattering processes. A precise and accurate determination of $\alpha_{s}(Q)$ and its dependence on the energy scale $Q$ is of vital importance for the current precision-frontier program at the Large Hadron Collider (LHC), where vast amounts of data are being gathered and push the boundary of precision measurements to the next level [1]. The upcoming high-luminosity phase of the LHC (HL-LHC) will further increase the volume of high-precision data, demanding a matching increase in both the detector capabilities and precision of theoretical calculations to properly validate relationships between fundamental parameters of the Standard Model (SM), constrain properties of the Higgs boson, and search for new physics. For instance, the running of $\alpha_{s}$ at the highest achieved energies may reveal the presence of new colored particles and influence the scale at which the SM couplings may unify. Variations in the observed $\alpha_{s}(M_{Z})$ at the scales of order of the $Z$ boson mass, $M_{Z}$, can affect projections for the range of energies at which the electroweak (EW) vacuum might shift from a metastable to a stable configuration.
@@
Hence, the QCD coupling and its estimated uncertainty play a pivotal dual role of characterizing the universal force of nature and, at the same time, serving as a precise theoretical input for many standard candle processes in ongoing experiments. In particular, top-quark pair and Higgs boson production cross sections are very sensitive to the assumed $\alpha_{s}(M_{Z})$. The associated uncertainty, which is correlated in hadron-collider predictions with the PDF uncertainty, can potentially exceed those due to the scale dependence and experimental errors [2, 3].
@@
The Particle Data Group (PDG) currently reports the uncertainty on $\alpha_{s}(M_{Z})$ of about 0.8%, resulting in a value of $\alpha_{s}(M_{Z})=0.118\pm 0.0009$  [4]. This result is derived from a compilation of lattice and non-lattice results, which independently provide compatible central values of $\alpha_{s}(M_{Z})$ and of its uncertainty. The corresponding relative uncertainty exceeds that achieved for the fundamental couplings of electroweak and gravitational interactions. The new twist is that the latest determinations of $\alpha_{s}$ from high-statistics experimental measurements and lattice QCD may individually quote a comparable and — perhaps in the future — a smaller uncertainty than the PDG world-average uncertainty of $\pm 0.0009$. Along these lines, the ATLAS collaboration measured $\alpha_{s}(M_{Z})=0.1183\pm 0.0009$ [5] using data of $Z$ boson transverse-momentum distributions at $\sqrt{s}=8$ TeV. The CMS collaboration simultaneously extracted $\alpha_{s}(M_{Z})$ and PDFs [6] from a global QCD analysis at NNLO using double-differential cross section measurements of inclusive jet production at $\sqrt{s}=$ 2.76, 7, 8, and 13 TeV combined with deep inelastic scattering (DIS) data from HERA. They obtained $\alpha_{s}(M_{Z})=0.1176^{+0.0014}_{-0.0016}$. Other recent relevant $\alpha_{s}$ determinations from the CMS and ATLAS collaborations can be found in Refs. [7, 8, 9, 5]. Several determinations from deep inelastic scattering (DIS) at $ep$ collider HERA find a somewhat lower $\alpha_{s}$ value than at the LHC, such as $\alpha_{s}(M_{Z})=0.1142\pm 0.0028$ determined by H1 collaboration in a simultaneous fit of $\alpha_{s}$ and proton PDFs [10]. Independently, lattice QCD determines $\alpha_{s}$ from several types of nonperturbative computations. The Flavor Lattice Averaging Group (FLAG) systematizes and cross-validates these determinations. It recently reports $\alpha_{s}(M_{Z})=0.1183\pm 0.0007$ [11].
@@
The studies of $\alpha_{s}$ thus enter a new phase when its precision may eventually be pushed to 0.1% across diverse observations in the next two decades [1], while in the meantime novel challenges will need to be confronted in both the individual $\alpha_{s}$ measurements and their combination. It is expected that, at this precision level, the growing complexity of involved systematic factors raises the chance for misestimations. When the misestimations go unnoticed – an unwanted but common occurrence encountered in complex studies across many disciplines [12] – the uncertainty tends to be underestimated; some measurements of the parameter are not replicated by its independent determinations. When replicability degrades, a PDG or another combination may need to inflate the stated uncertainties of measurements in order to make them statistically compatible. As an illustration, the Committee on Data of the International Science Council inflates the errors of highly precise measurements of Newton’s gravitational constant $G$ by a factor of 3.9 in order to meaningfully combine them [13]. Within its small span, the combined uncertainty band on $G$ does not overlap significantly with those of several most precise measurements, meaning it does not replicate them. This indicates a general trend that the measurements of fundamental constants eventually experience the tradeoff between the precision and replicability, when raising the former depletes the confidence in the latter. How far can the precision measurements of $\alpha_{s}$ go while remaining replicable? What can be done to control the uncertainty estimates and postpone the onset of non-replicability?
@@
This article will critically explore these questions in the context of the precision determination of the $\alpha_{s}(M_{Z})$ from the new CTEQ Tung Et Al. (CTEQ-TEA) global QCD analysis of nucleon scattering data, named “CT25” [14]. The primary goal of the global QCD analysis is to determine parametrizations of parton distribution functions (PDFs), as well as $\alpha_{s}$ and other QCD parameters, from a simultaneous fit of a large, diverse collection of experimental measurements. The methodology of the global analysis offers the most complete view of the correlated constraints on $\alpha_{s}$, heavy-quark masses, and the PDFs arising from a variety of scattering experiments with different systematic errors, with simultaneous dependence on all of them. This completeness gives the global analysis an edge over the standalone $\alpha_{s}$ determinations from the nucleon data, which typically neglect at least some of the correlation with the PDFs or other experiments. As such, and given the growing dominance of hadron scattering processes in experimental $\alpha_{s}$ determinations, the global analysis of PDFs is well suited for tests of uncertainty quantification and replicability in the $\alpha_{s}$ studies. Many combined analyses of the PDFs and $\alpha_{s}$ have been performed by our CTEQ-TEA group [15, 16] and other groups [17, 18, 19, 20, 21, 22] over the years. The strong coupling $\alpha_{s}$ has recently been extracted from global QCD analyses, employing an approximate PDF evolution at the next-to-next-to-next-to-leading order (aN3LO) [23] in QCD by the MSHT group, and from an analysis using an aN3LO QCD evolution combined with the photon PDF at the next-to-leading order (NLO) in QED [24] by the NNPDF collaboration.
@@
The CT25 global analysis determines $\alpha_{s}$ at the next-to-next-to-leading order (NNLO) in QCD by employing an improved methodology as well as an extended global data set in which precision LHC Run-2 measurements with high integrated luminosity augment those included in the CT18 study [16]. The novel feature of our study is that we determine the central value and uncertainty of $\alpha_{s}$ by several methods to obviate the dependence on the correlation with PDFs, selection of experiments, and systematic uncertainties in the experiments.
@@
The rest of this manuscript is organized as follows. In Sec. II, we briefly describe the methodology and advancements in the CT25 global PDF analysis. We summarize the new data sets included in the CT25 baseline, the definition of the figure-of-merit function $\chi^{2}$, and the implementation of correlated systematic errors provided by the experiments. Section III.1 explores the impact of the choice of PDF parametrizations on the $\alpha_{s}$ determination. We also review how $\alpha_{s}(M_{Z})$ can be determined from the $\chi^{2}$ scan over $\alpha_{s}(M_{Z})$, and which criteria are applicable to determine the $\alpha_{s}(M_{Z})$ error with this methodology.
@@
Sec. III.2 critically examines the impact of the implementation of experimental systematic uncertainties on the extracted value of $\alpha_{s}(M_{Z})$, and it discusses the handling of some observed disagreements among data sets in the baseline. Section IV is dedicated to the exploration of uncertainty prescriptions (tolerances) according to the commonly used global and dynamical implementations. For the latter definition, we examine how the user-specified clustering of published data sets modifies the quantitative error estimates. We point out that either the dynamical tolerance in the Hessian method or $K$-folding in the Monte-Carlo method may produce uncertainties that depend on the user-chosen clustering of data sets. We introduce the concept of data clustering safety that must be addressed in such situations.
@@
Sec. V describes Bayesian statistical models utilized for the combination of the results for $\alpha_{s}(M_{Z})$ and its uncertainties, while Sec. VI obtains and compares the uncertainties using a variety of methods. This Section quotes our final $\alpha_{s}$ estimate. Sec. VII contains our conclusions. Technical details of several discussions are relegated to five Appendices. The default scale $M_{Z}$ is to be assumed when the scale of $\alpha_{s}$ is omitted.
II Relevant aspects of the CT25 global analysis
II.1 Selection of data sets
As in many general-purpose CTEQ-TEA fits, the CT25 NNLO global analysis has taken place over several years and now culminates with the release of the official CT25 PDF ensembles that update the previous generation of the PDFs [16] consisting of the default CT18 NNLO error set and three alternative, yet equally acceptable PDF ensembles, CT18Z, A, and X. A significant part of this effort went into an exploration of a large number of new candidate measurements for the extension of the CT18 baseline data set, focusing primarily on high-precision production of lepton pairs, hadronic jets, and top quark-antiquark pairs in the LHC Runs 1 and 2. We have documented these preliminary explorations in three publications [26, 27, 28]. Based on these studies, we selected a number of precise and mutually consistent data sets that are included in the final CT25 release. Several data sets included in CT18 have been updated or removed in the final CT25 fit, for the reasons that will be explained in the text.
The newly added LHC data sets in CT25 are listed in the upper part of Table 1 and reviewed in Appendix A. As its data baseline, CT25 started with the CT18A [16] global data set, which already included a combination of measurements in deep-inelastic scattering and production of lepton pairs, hadronic jets, and top-quark pairs at both fixed-target and collider energies. Notably, the CT18A baseline included a precision data set on $W$ and $Z$ production in the ATLAS experiment at 7 TeV [29], which offered the first evidence for the enhanced strangeness PDF at $x\approx 0.02$, in some disagreement with charged-current DIS experiments. More recent LHC measurements included in the CT25 fit are consistent with this strangeness enhancement, albeit with somewhat reduced magnitude.
The inclusion of the new LHC measurements, added on top of the CT18A baseline without replacing or discarding any of the previously fitted experiments, increased the number of data points, $N_{\textrm{pt}}$, by 769, the majority of which were from inclusive jet production and Drell-Yan triple-differential $Z$-boson production at ATLAS, cf. Table 1. This version of the CT25 fit, which we will call “CT25prel”, has the largest $N_{\textrm{pt}}=4450$ among the various CT25 candidate fits. This data set is not unique, however: other well-motivated data combinations are possible, as was also the case for CT18. When all data sets are retained in the CT25 baseline, we observe a reasonable, although not ideal, agreement among the global data, with significant pulls on the PDFs from the newly-fitted data. The most precise new experiments in the CT25 fits tend to exhibit an increase in $\chi^{2}/N_{\textrm{pt}}$ in a pattern similar to that observed in the CT18 global analysis. The new Drell-Yan and $t\bar{t}$ production data consistently prefer a softer gluon at intermediate-to-large $x$ ($0.05\leq x\leq 0.7$), while inclusive jet production data pull the gluon in the opposite direction in the large-$x$ range. Overall, the cumulative effect of these measurements, when they are combined, amount to having a softer gluon at intermediate-to-large $x$, as compared to CT18A.
To investigate the range of valid PDF solutions in this new setting, we explored other related, also acceptable, branches of the CT25prel fit. The alternative choices made in these fits reflect the complexities present within the global fitting framework, many of which were highlighted in the preceding CT18 publication [16], as well as in a review [25]. [^1] In the context of CT18, we opted to publish four PDF ensembles to quantify some of these possibilities. In the end, we obtained the default CT25 fit, or just “CT25”, as follows. Compared to CT18A and CT25prel, we replaced the H1 measurements of charm and bottom semi-inclusive structure functions [30, 31] by their HERA combination [32], which moderately increased the corresponding $\chi^{2}$ without substantially modifying the PDFs. We added the E906/SeaQuest measurement of the $\sigma_{pd}/(2\sigma_{pp})$ ratio in fixed-target Drell-Yan pair production [33]. We removed four data sets (336 points) on inclusive charged-current DIS on an iron target from the CDHSW [34] and CCFR [35, 36] collaborations, as they are in some tension with each other and increase dependence of CT25 results on the nuclear modifications of proton PDFs. [^2] The default CT25 fit nevertheless retains CCFR and NuTeV data sets on dimuon production in CC DIS on the isoscalar target, as those provide unique constraints on the strangeness PDF at $x>0.05$. In both CT18 and CT25 fits, we apply a correction from [37] to approximate nuclear modifications in such experiments. The weakly sensitive determination of the longitudinal structure function $F_{L}(x_{B},Q^{2})$ by the H1 collaboration [38] was also omitted without a noticeable impact on the PDFs. Finally, we replaced the CMS jet production datasets at 7, 8, and 13 TeV by their updated versions [6] that incorporate improved evaluations of systematic uncertainties and include statistical correlation matrices.
Table 1: The upper part of the table lists the newly added data sets in the CT25 analysis, contributed by measurements of lepton-pair, top-quark-pair, and single-inclusive jet production from the LHC Runs 1 and 2. For these experiments, we report the $\chi^{2}/N_{\textrm{pt}}$ for the default CT25 and prior CT18A PDF ensembles, assuming $\alpha_{s}(M_{Z})=0.118$. Section II.2 provides the definition of $\chi^{2}$. The errors are asymmetric PDF uncertainties computed using a master formula in Ref. [39] and the error PDFs defined at 68% CL, according to the two-tier PDF tolerance employed in CT18. The lower part of the table shows the total numbers of points, $\chi^{2}/N_{\textrm{pt}}$, and approximate preferred $\alpha_{s}$ in the default and preliminary CT25 fits described in the text, as well as in the CT18A fit.
New and updated LHC data sets in the CT25 NNLO analysis
@@ II.1 Selection of data sets
Inclusive jet production
553 ATLAS 8 TeV jets [46] 171 $1.60^{+0.10}_{-0.06}$ ( $1.57^{+0.12}_{-0.07}$ )
554 ATLAS 13 TeV jets [47] 177 $1.32^{+0.06}_{-0.05}$ ( $1.26^{+0.09}_{-0.03}$ )
555 CMS 13 TeV jets [48, 6] 78 $1.13^{+0.11}_{-0.04}$ ( $1.29^{+0.23}_{-0.17}$ )
556 CMS 7 TeV jets [49, 50, 51, 6] 118 $0.74^{+0.02}_{-0.03}$ ( $0.7^{+0.12}_{-0.03}$ )
557 CMS 8 TeV jets [52, 6] 164 $1.17^{+0.12}_{-0.06}$ ( $1.16^{+0.18}_{-0.06}$ )
@@
CT25prel, with augmented CT18A data set 4450 $1.20\pm 0.010$ $\approx 0.1177$
@@
Turning back to Table 1, the last column in its upper portion reports the $\chi^{2}/N_{\rm pt}$ values for each experiment using the PDFs from the default CT25 NNLO and from the CT18A NNLO global fits (with the latter in parentheses). The errors express the corresponding asymmetric PDF uncertainties that are evaluated at the 68% credibility level (CL). [^3] The PDF fits base themselves on the Bayesian paradigm, and hence the interpretation of their errors on the PDFs and $\alpha_{s}$ as “credibility intervals” is appropriate. That said, these errors are also used to predict the frequency of future measurements using the determined parameters, in which case the users may view them as “confidence intervals”. Fitting the new experiments in the CT25 framework generally improves their $\chi^{2}$.
@@
The lower portion of Table 1 compares the total numbers of points and $\chi^{2}/N_{\textrm{pt}}$ for the default CT25, CT25prel, and their CT18A predecessor. The last column of the lower part also shows the approximate preferred value of $\alpha_{s}$ in these three fits, which is further adjusted by modifications discussed in the rest of the article.
@@
Upon the addition of the LHC data sets in CT25prel, the reference $\alpha_{s}$ increased from a range 0.1164-0.1169 in the CT18 series to $\approx 0.118$. To a large extent, this change is due to the preference for a softer gluon at the intermediate-to-large $x$, which is compensated by some increase in $\alpha_{s}$.
@@
In the default CT25 fit, the reference $\alpha_{s}$ is further increased to $\approx 0.1185$ primarily because of the removal of the CDHSW inclusive charged-current DIS data set [34] on iron. The exclusion of the counterpart CCFR measurement had no appreciable effect, as it is known to be largely consistent with $\alpha_{s}(M_{Z})\approx 0.118$[57, 58, 59, 60]. In the CT18 publication, we emphasized that the CDHSW data set is generally fitted well, yet its $Q$ dependence differs from that of an analogous CCFR observation [61, 62].
The Lagrange Multiplier scans in the context of CT18 (see its Fig. 21 in Ref. [16]) indicated that CDHSW prefers a far larger gluon $g(x,Q)$ at $x>0.1$ than the other experiments. This preference, in turn, suppresses the preferred $\alpha_{s}$. For this reason, the two CDHSW data sets were included in CT18/CT18A/CT18X, but not in CT18Z. Aside from the impact on the fitted value of $\alpha_{s}$, the removal of the inclusive neutrino-iron DIS data sets from the CT25prel PDF fit slightly (well within PDF uncertainties) reduced $g$ and increased $u$ and $d$ PDFs at $x>0.1$. Overall, these CDHSW and CCFR experiments do not manifestly disagree with the rest of the data.
@@
Our general observation is then that the described revisions in the CT25 data set, as compared to the CT18A data set, tend to elevate the preferred $\alpha_{s}$ to above 0.118 for the default settings of the global fit. We further emphasize that this trend may change considerably under reasonable modifications in the fit’s settings, reflecting in part inconsistencies (tensions) among the input data sets, and in part the modeling of experimental and theoretical systematic uncertainties.
Many CPU-intensive studies were carried out with the CT25prel data set, and were not repeated for the (slightly different) CT25 data set. In the rest of this article, we will present studies with either CT25prel or with CT25, without loss of generality.